1/5+3yequals5y-17/25
Answer by unlockmath(1688)

Hello,
Let's rewrite this 1/5+3yequals5y-17/25 as:
1/5+3y=5y-17/25
Now multiply by 25 to get:
5+75y=125y-17
Subtract 75 and add 17 to get:
50y=22
Divide by 50 to get:
y=22/50 or 11/25
Make sense?
RJ
www.math-unlock.com